OpenStudy (anonymous):

what does y equal in this equation? -4x+7y=-28

OpenStudy (anonymous):

-4x+7y = -28 +4x +4x 7y = -28+4x then divide both sides by 7 which equals y = (4x-28)/(7)

OpenStudy (anonymous):

Thank you!!(: